OpenStudy (anonymous):

What is the solution of the equation 7/9x+8=5x

OpenStudy (anonymous):

\[\frac{ 7 }{ 9 }x + 8 = 5x\]

OpenStudy (anonymous):

Let's solve your equation step-by-step. 7/9x+8=5x Step 1: Subtract 5x from both sides. 7/9x+8−5x=5x−5x −38/9x+8=0 Step 2: Subtract 8 from both sides. −38/9x+8−8=0−8 −38/9x=−8 Step 3: Multiply both sides by 9/(-38). (9/−38)*(−38/9x)=(9−38)*(−8) x=36/19

OpenStudy (anonymous):

Thank youuu

OpenStudy (anonymous):

yw